If the economy booms, RTF, Inc., stock is expected to return 12 percent. If the economy goes into a recessionary period, then RTF is expected to only return 4 percent. The probability of a boom is 76 percent while the probability of a recession is 24 percent. What is the variance of the returns on RTF, Inc., stock? Multiple Choice .001167 .000887 .034167 .050400
Answer #1

E(r)=(0.76*0.12)+(0.24*0.04)

=0.0912+0.0096

=0.1008


variance=0.76* (0.12-0.1008)^2 +0.24* (0.04-0.1008)^2

= 0.0002801664+0.0008871936

=0.001167